OpenStudy (anonymous):

quadratic equation.....(y+6)squared=1

OpenStudy (anonymous):

t=y+6 t^2=1 \[t=\pm1\] \[y+6=\pm1\]

OpenStudy (anonymous):

\[y=\pm1-6\]

OpenStudy (anonymous):

y=-7 or y=-5